OpenStudy (anonymous):

What is equal to √200?

OpenStudy (anonymous):

\[\sqrt{200}=\sqrt{2}\sqrt{100}\]

OpenStudy (anonymous):

10√2

OpenStudy (anonymous):

Or 14.1421

OpenStudy (anonymous):

thanx